OpenStudy (anonymous):

Which set of angles shows an acute triangle? twenty degrees, sixty degrees, one hundred degrees fifteen degrees, one hundred fifty degrees, fifteen degrees thirty-five degrees, ninety degrees, fifty-five degrees fifty degrees, sixty degrees, seventy degrees

OpenStudy (mertsj):

Acute triangles have 3 acute angles.

OpenStudy (anonymous):

Im Lost even more what?

OpenStudy (anonymous):

Acute means less than 90 , right?

OpenStudy (anonymous):

Let us take an angle, which is acute : \(\theta\) Then : \(0 < \theta < 90 \) Can you tell me now that what is the sum of all the angles of a triangle?

OpenStudy (anonymous):

No?

OpenStudy (anonymous):

Im stuiped?

OpenStudy (mertsj):

An acute angle is less than 90 degrees. An acute triangle has three acute angles. Which answer has three acute angles?

OpenStudy (anonymous):



OpenStudy (anonymous):

Note this formula : Sum of all angles of a polygon = (n-2) * 180 where n = number of sides Here, in a triangle there are three sides, so n = 3 Put this in the formula : Sum of all angles of a polygon ( here , it is triangle) = (3-2) * 180 degrees

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

Oh The last One!

OpenStudy (mertsj):

Bingo.

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

Kitty, concentrate on mertsj answer. Please answer his question.

OpenStudy (anonymous):

Do what?

OpenStudy (anonymous):

Good. That's right. Also , \(\huge{\color{blue}{\mathsf{Welcome }} \space \color{orange}{\mathsf{To}} \space \color{gold}{\mathsf{OpenStudy!!!}}} \)

OpenStudy (anonymous):

Do nothing now :)

OpenStudy (anonymous):

Haha No I still do it its just when Im stuck I need help But thank you